OpenStudy (anonymous):

I have several problems similar to : find tan theta if (-5,12) is a point on the terminal side of theta

OpenStudy (anonymous):

x = -5 y = -12 r = 13 sin θ = y/r = -12/13 cos θ = x/r = -5/13 tan θ = y/x = 12/5

OpenStudy (anonymous):

where did the 13 come from?

OpenStudy (anonymous):

to find the hypotenuse you take 5^2 + 12^2 and then you will get 169 which is 13^2 so I get that I have all 3 sides

OpenStudy (anonymous):

do you needmore help

OpenStudy (amistre64):

tan(x) = y x = tan-1(y)

OpenStudy (amistre64):

tan(t) = -12/5 t = tan-1(-12/5) thats the inverse tan function

OpenStudy (amistre64):

what it amounts to i dont know; have to calulator it for an exact theta(t)

OpenStudy (anonymous):

good